Hi,
Could you explain how I would do this word problem? Thanks!
The perimeter of a triangle is 24cm. One side is 2cm longer than the shortest side. The longest side is 2cm less than twice the length of the shortest side. Find the length of each side of the triangle.

Answer by stanbon(75887)

The perimeter of a triangle is 24cm. One side is 2cm longer than the shortest side. The longest side is 2cm less than twice the length of the shortest side. Find the length of each side of the triangle.
Let the shortest side have length "x" cm.
A second side is "x+2" cm.
The longest side is "2x-2" cm.
EQUATION:
The sum of the sides is 24 cm
x+x+2+2x-2=24
4x=24
x=6 cm (length of shortest side)
x+2=8 cm (length of second side)
2x-2=10 cm (length of longest side)
Cheers,
stan H.